OpenStudy (anonymous):

a right triangle has legs of 8 units and 10 units use coordinate plane to love for the hypotenuse

OpenStudy (anonymous):

 To solve use the Pythagorean Theorem: a^2+b^2=c^2 Where a and b are the legs. But I guess that doesn't really use the coordinate plane

OpenStudy (anonymous):

A can be any number , it cab be 8 or 10 you chose

OpenStudy (anonymous):

then having the a=8 and 10=b you have c for you to solve

OpenStudy (anonymous):

so next you square 8^2 and 10^2 8=64, 10=100 it give's you 164

OpenStudy (anonymous):

did you get it?